# Complex crystallographic reflection groups and Seiberg–Witten integrable systems: rank 1 case

### v2

@@ 1 Introduction
In our previous paper [2], we proposed to study the so-called crystallographic elliptic Calogero–Moser systems as a potential source of Seiberg–Witten integrable systems. Recall that, according to [1], each complex crystallographic reflection group has an associated family of elliptic Calogero–Moser systems constructed using the theory of elliptic Cherednik algebras. We expect that many of these systems can be viewed as Seiberg–Witten integrable systems for certain superconformal field theories (SCFTs). In [2], this was partly confirmed for the Inozemtsev system, a $BC_{n}$-version of the elliptic Calogero–Moser system associated to the group $W=\mathbb{Z}_{2}\wr S_{n}$.
In the present paper we turn our attention to groups (and SCFTs) of rank one. Geometrically, this means considering elliptic curves $\mathcal{E}$ with $\mathbb{Z}_{m}$-symmetries, $m=2,3,4,6$, and Poisson deformations of the orbifolds $T^{*}\mathcal{E}/\mathbb{Z}_{m}$. Such deformations can be constructed using the classical version of the Cherednik algebra, $H_{c}$, and its spherical subalgebra, $B_{c}$. Here $c$ denotes the set of deformation parameters whose number equals $4$, $6$, $7$ and $8$ for $m=2,3,4$ and $6$, respectively. The integrable system appears as the (Poisson-commutative) algebra of the global sections of $B_{c}$. In our case, it is described by a single hamiltonian function $h$ on $T^{*}\mathcal{E}$, which can also be viewed as a function on $T^{*}\mathcal{E}/\mathbb{Z}_{m}$ due to its symmetry.
@@
Turning to the supersymmetric quantum filed theories, one would like to associate the geometric picture above with some rank 1 4d $\mathcal{N}=2$ SCFTs. For the rich landscape and classification of such theories, see [9, 10, 11, 12]. As it turns out, our $m=2$ case corresponds to the well-known ${\rm SU}(2)$ superconformal gauge theory with 4 flavors which possesses a $D_{4}$ flavor symmetry [7]. Meanwhile, the $m=3,4,6$ cases correspond to Minahan-Nemeschansky (MN) theories [3, 4] that possess the exceptional flavor symmetry of types $E_{6}$, $E_{7}$, $E_{8}$ (note that these theories lack conventional Lagrangian descriptions). This new perspective allows us to obtain the elliptic fibration and the Seiberg–Witten differential for the rank 1 MN theories in a systematic fashion. Notably, the mass parameters of those rank 1 SCFTs receive a transparent geometric interpretation, at the same time being directly linked to the deformation parameters of the elliptic Cherednik algebras. As a byproduct, we find a natural quantisation of the Seiberg–Witten curves of the rank 1 Minahan–Nemeschansky theories. It is worth mentioning that the elliptic fibrations in Weierstrass form have already been established for these theories [6, 3, 4], but they seem less suitable for quantisation.
@@ 2 Complex crystallographic groups and Cherednik algebras in rank one
If $X$ is a complex manifold with an action of a finite group $W$, that action naturally extends to the sheaf $\mathcal{D}[X]$ of regular differential operators on $X$. Therefore, one may consider $\mathcal{D}[X]\rtimes W$ and $\mathcal{D}[X]^{W}$ as sheaves of algebras over $X/W$. In such situation, Etingof constructs in [14] the global Cherednik algebra $H_{c}(X,W)$ and its spherical subalgebra $B_{c}(X,W)$ as certain (in fact, universal) deformations of $\mathcal{D}[X]\rtimes W$ and $\mathcal{D}[X]^{W}$, respectively. A special case of interest is when $X={\mathbb{C}}^{n}/\Gamma$ is a complex torus, and $W\subset\mathrm{GL}_{n}({\mathbb{C}})$ is a complex reflection group preserving the lattice $\Gamma$, thus acting on $X$. The semi-direct product $G=\Gamma\rtimes W$ is an example of a complex crystallographic group; all such groups are classified in [15] [^1] In general, $G$ may not be a semidirect product of $W$ and $\Gamma$. Also, one may consider a more general situation than in [15] so that $W$ is generated by reflections but $G$ may not be (that is how, for example, extended Weyl groups arise).. In this case $H_{c}(X,W)$ is referred to as the elliptic Cherednik algebra. The significance of $X$ being a complex torus is that in that case, according to [1], the spherical subalgebra $B_{c}(X,W)$ has a commutative subalgebra of dimension $n$. This defines a family of integrable systems on $X$, called crystallographic elliptic Calogero–Moser systems. In this paper we look into the simplest cases, namely, complex crystallographic groups and elliptic Cherednik algebras of rank 1, corresponding to elliptic curves with symmetries.
@@ 2.1 Elliptic curves with symmetries
and write $\mathbb{Z}_{m}:=\{1,s,\dots s^{m-1}\}$ for the multiplicative group of the $m$-th roots of unity, acting on $\mathcal{E}$. The generic $\mathcal{E}$ corresponds to $m=2$. The point $q=0$ is always fixed by $\mathbb{Z}_{m}$; other fixed points and their stabiliser groups are given in the table 1. These are also shown in figure 1.
$\begin{array}[]{c|c|c }\hline\cr\qquad m&\qquad\text{fixed points$\neq 0$}&\qquad\text{stabilisers}\\ \hline\cr 2&\omega_{1,2,3}&\mathbb{Z}_{2}\\ \hline\cr 3&\eta_{1,2}&\mathbb{Z}_{3}\\ \hline\cr 4&\begin{array}[]{c}\omega_{1,2}\\ \omega_{3}\end{array}&\begin{array}[]{c}\mathbb{Z}_{2}\\ \mathbb{Z}_{4}\end{array}\\ \hline\cr 6&\begin{array}[]{c}\omega_{1,2,3}\\ \eta_{1,2}\end{array}&\begin{array}[]{c}\mathbb{Z}_{2}\\ \mathbb{Z}_{3}\end{array}\\ \hline\cr\end{array}$
@@
Let $\sigma(q)=\sigma(q|2\omega_{1},2\omega_{2})$, $\zeta(q)=\zeta(q|2\omega_{1},2\omega_{2})$, $\wp(q)=\wp(q|2\omega_{1},2\omega_{2})$ be the Weierstrass $\sigma$, $\zeta$ and Weierstrass-$\wp$ functions associated with $\Gamma$ and $\mathcal{E}$. Since $\omega\Gamma=\Gamma$, we have
@@
$\begin{array}[]{c | c c c c}\hline\cr m&2&3&4&6\\ \hline\cr\omega_{2}/\omega_{1}&\text{any}&e^{{\pi i}/{3}}&i&e^{{\pi i}/{3}}\\ \hline\cr u&\wp(q)&\frac{1}{2}\wp^{\prime}(q)&\wp^{2}(q)&\wp^{3}(q)\\ \hline\cr v&\frac{1}{2}\wp^{\prime}(q)&\wp(q)&\frac{1}{2}\wp^{\prime}(q)&\frac{1}{2}\wp(q)\wp^{\prime}(q)\\ \hline\cr P_{m}(u)&u^{3}-\frac{1}{4}g_{2}u-\frac{1}{4}g_{3}&u^{2}+\frac{1}{4}g_{3}&u(u-\frac{1}{4}g_{2})^{2}&u^{2}(u-\frac{1}{4}g_{3})^{3}\\ \hline\cr\end{array}$
@@ 2.2 Rational Cherednik algebra for $W=\mathbb{Z}_{m}$
Then one finds that under $\theta$, the elements $e\,q^{m}\,e$, $e\,y^{m}\,e$, and $e\,qy\,e$ are mapped to $u,w$, and $v-\mu_{0}$, respectively. We denote by $B_{\hbar,c}$ the spherical subalgebra in the Dunkl representation:
@@
$[v,u]=\hbar mu\,,\quad[w,v]=\hbar mw\,,\quad uw=P_{\hbar}(v)\,,\quad wu=P_{\hbar}(v+\hbar m)\,,$ (2.20)
@@ 2.3 Elliptic Cherednik algebra of rank one
We now proceed to define the elliptic version of $H_{\hbar,c}$, following the general framework of [14]. Let $\mathcal{E}=\mathbb{C}/\Gamma$ be an elliptic curve with the symmetry group $\mathbb{Z}_{m}$, $m\in\{2,3,4,6\}$. The elliptic Cherednik algebra $H_{\hbar,c}(\mathcal{E})$ depends on a set of parameters chosen as follows. To every $x_{i}\in\mathcal{E}$ and $l=1,\dots,m-1$ such that $s^{l}(x_{i})=x_{i}$, we assign a parameter $c_{l}(x_{i})$, with an additional requirement that $c_{l}(x_{i})=c_{l}(x_{j})$ whenever $x_{j}=wx_{i}$ for some $w\in W$. From Fig. 1 we observe that this amounts to $4$ parameters in the $m=2$ case, and $6$, $7$ or $8$ parameters when $m=3$, $4$ or $6$, respectively. We write $c=(c_{l}(x_{i}))$ for the set of parameters. It will be convenient to extend the set $c$ by setting $c_{l}(x_{i})=0$ if $s^{l}(x_{i})\neq x_{i}$. Later it will also be convenient to use the following combinations:
@@
This defines the sheaf $H_{\hbar,c}(\mathcal{E})$ of elliptic Cherednik algebras on $\mathcal{E}$. The sheaf of spherical subalgebras $eH_{\hbar,c}(\mathcal{E})e$ is obtained by replacing local sections $a\in\mathcal{D}_{\mathcal{E}}\rtimes\mathbb{Z}_{m}$ by $eae$. Again, these local sections can be realised as differential operators using the map (2.16). This produces a sheaf of algebras $B_{\hbar,c}(\mathcal{E}):=\theta\left(eH_{\hbar,c}(\mathcal{E})e\right)\subset\mathcal{D}_{\mathcal{E}}^{\mathbb{Z}_{m}}$.
@@ 2.5 Hamiltonians
To prove this, we notice that any global section $\widehat{h}$ restricts onto $\mathcal{E}_{0}=\mathcal{E}\setminus\{0\}$, and so it must be a combinations of elements $w_{i}$, $v^{(i)}_{j}$. For $\widehat{h}$ to have degree $m$ in $\widehat{p}$, it must be obtained from $w_{m}$ by adding a finite linear combination of $v^{(i)}_{j}$ with $10\leq j\leq m-$ 1. On the other hand, near $q=0$ each global section must belong to the (completion) of the local rational spherical Cherednik subalgebra, generated by $\mathbb{C}[[u]]$, $v$ and $w$ as given in (2.18). Since $u,v$ are regular at $q=0$, we must have
@@
Comparing this with the previous formula, we conclude that the only allowed terms in $\widehat{h}$ are $v^{(i)}_{j}$ with $j=0,\dots,m-2$ and $i+j=m-2$, thus proving (2.40).
@@ 2.6 Explicit formulas
Let us write the hamiltonians explicitly, based on the formula (2.40). We use the notation $\omega_{1,2,3}$ and $\eta_{1,2}$ for the fixed points, as in (1). The parameters $\alpha_{i}$ below are determined through (2.44).
@@ 3.1 Classical dynamics
defines a non-vanishing holomorphic $1$-form on $\Sigma$. Hence, the $\mathbb{Z}_{m}$-quotient of the fibration (3.2) defines an elliptic fibration on $T^{*}\mathcal{E}/\mathbb{Z}_{m}$, and the dynamics (3.1) becomes linear along the fibers.
@@ 3.2 Elliptic Dunkl operators
Let us fix some notation. For an abelian variety $X$ with an action of a finite group $W$, the Dunkl operators [19] depend on an auxiliary variable $\alpha\in X^{\vee}=\mathrm{Pic}_{0}(X)$. We will be dealing with the case $X=\mathcal{E}=\mathbb{C}/\Gamma$ and $W=\mathbb{Z}_{m}$, in which case we identify $X^{\vee}\simeq X$ so that $\alpha\in\mathcal{E}$. Below $\sigma$, $\zeta$, Weierstrass-$\wp$ stand for the Weierstrass functions associated to the lattice $\Gamma=2\omega_{1}\mathbb{Z}+2\omega_{2}\mathbb{Z}$, and we write $\varphi(x,z)$ for the following combination:
@@
Recall that for each fixed point $x_{i}\in\mathcal{E}$ we have parameters $c_{l}(x_{i})$, $11\leq l\leq m-$ 1, with $c_{l}(x_{i})=0$ unless $s^{l}(x_{i})=x_{i}$. The fixed points of $s^{l}$ are identified with the cosets
@@
$v_{l}(x,z)$ =$=\sum_{\{x_{i}\}}c_{l}(x_{i})e^{-\eta(\Omega_{l}x_{i})z}\varphi(x-x_{i},\Omega_{-l}z)\,,\qquad l\in\mathbb{Z}_{m}\setminus\{0\}\,,$ (3.8)
@@
[…]

### v3

@@ 1 Introduction
In our previous paper [2], we proposed to study the so-called crystallographic elliptic Calogero–Moser systems as a potential source of Seiberg–Witten integrable systems. Recall that, according to [1], each complex crystallographic reflection group has an associated family of elliptic Calogero–Moser systems constructed using the theory of elliptic Cherednik algebras. We expect that many of these systems can be viewed as Seiberg–Witten integrable systems for certain superconformal field theories (SCFTs). In [2], this was partly confirmed for the Inozemtsev system, a version of the elliptic Calogero–Moser system associated with the non-reduced root system of type $BC_{n}$, or with the complex crystallographic group $G(1,2,n)]^{\alpha}_{1}$ in Popov’s classification [8, Theorem 2.6.1].
In the present paper we turn our attention to groups (and SCFTs) of rank one. Geometrically, this means considering elliptic curves $\mathcal{E}$ with $\mathbb{Z}_{m}$-symmetries, $m=2,3,4,6$, and Poisson deformations of the orbifolds $T^{*}\mathcal{E}/\mathbb{Z}_{m}$. Such deformations can be constructed using the classical version of the Cherednik algebra, $H_{c}$, and its spherical subalgebra, $B_{c}$. Here $c$ denotes the set of deformation parameters whose number equals $4$, 66, 77 and $8$ for $m=2,3,4$ and $6$, respectively. The integrable system appears as the (Poisson-commutative) algebra of the global sections of $B_{c}$. In our case, it is described by a single hamiltonian function $h$ on $T^{*}\mathcal{E}$, which can also be viewed as a function on $T^{*}\mathcal{E}/\mathbb{Z}_{m}$ due to its symmetry.
@@
Turning to the supersymmetric quantum field theories, one would like to associate the geometric picture above with some rank 1 4d $\mathcal{N}=2$ SCFTs. For the rich landscape and classification of such theories, see [10, 11, 12, 13]. As it turns out, our $m=2$ case corresponds to the well-known ${\rm SU}(2)$ superconformal gauge theory with 4 flavors which possesses a $D_{4}$ flavor symmetry [7]. Meanwhile, the $m=3,4,6$ cases correspond to Minahan-Nemeschansky (MN) theories [3, 4] that possess the exceptional flavor symmetry of types $E_{6}$, $E_{7}$, $E_{8}$ (note that these theories lack conventional Lagrangian descriptions). This new perspective allows us to obtain the elliptic fibration and the Seiberg–Witten differential for the rank 1 MN theories in a systematic fashion. Notably, the mass parameters of those rank 1 SCFTs receive a transparent geometric interpretation, at the same time being directly linked to the deformation parameters of the elliptic Cherednik algebras. As a byproduct, we find a natural quantisation of the Seiberg–Witten curves of the rank 1 Minahan–Nemeschansky theories. It is worth mentioning that the elliptic fibrations in Weierstrass form have already been established for these theories [6, 3, 4], but they seem less suitable for quantisation.
@@ 2 Complex crystallographic groups and Cherednik algebras in rank one
If $X$ is a complex manifold with an action of a finite group $W$, that action naturally extends to the sheaf $\mathcal{D}[X]$ of regular differential operators on $X$. Therefore, one may consider the crossed product $\mathcal{D}[X]\rtimes W$ and the subsheaf of $W$-invariants $\mathcal{D}[X]^{W}$, both regarded as sheaves of algebras over $X/W$. In such situation, Etingof constructs in [15] the global Cherednik algebra $H_{c}(X,W)$ and its spherical subalgebra $B_{c}(X,W)$ as certain (in fact, universal) deformations of $\mathcal{D}[X]\rtimes W$ and $\mathcal{D}[X]^{W}$, respectively. A special case of interest is when $X={\mathbb{C}}^{n}/\Gamma$ is a complex torus, and $W\subset\mathrm{GL}_{n}({\mathbb{C}})$ is a complex reflection group preserving the lattice $\Gamma$, thus acting on $X$. The semi-direct product $G=\Gamma\rtimes W$ is an example of a complex crystallographic reflection group, namely a discrete subgroup of the complex affine group of $\mathbb{C}^{n}$ generated by (affine) reflections, whose linear part $W$ is a finite complex reflection group and whose translation lattice $\Gamma$ is of full rank $2n$ in $\mathbb{C}^{n}$; all such groups are classified in [8] [^1] In general, $G$ may not be a semidirect product of $W$ and $\Gamma$. Also, one may consider a more general situation than in [8] so that $W$ is generated by reflections but $G$ may not be (that is how, for example, extended affine Weyl groups arise).. In this case $H_{c}(X,W)$ is referred to as the elliptic Cherednik algebra. The significance of $X$ being a complex torus is that in that case, according to [1], the spherical subalgebra $B_{c}(X,W)$ has a commutative subalgebra of dimension $n$. This defines a family of integrable systems on $X$, called crystallographic elliptic Calogero–Moser systems. In this paper we look into the simplest cases, namely, complex crystallographic groups and elliptic Cherednik algebras of rank 1, corresponding to elliptic curves with symmetries.
@@ 2.1 Elliptic curves with symmetries
and write $\mathbb{Z}_{m}:=\{1,s,\dots s^{m-1}\}$ for the multiplicative group of the $m$-th roots of unity, acting on $\mathcal{E}$. The generic $\mathcal{E}$ corresponds to $m=2$. The point $q=0$ is always fixed by $\mathbb{Z}_{m}$; other fixed points and their stabiliser groups are given in the table 1. These are also shown in Fig. 1.
[…]
@@
Let $\sigma(q)=\sigma(q|2\omega_{1},2\omega_{2})$, $\zeta(q)=\zeta(q|2\omega_{1},2\omega_{2})$, $\wp(q)=\wp(q|2\omega_{1},2\omega_{2})$ be the Weierstrass $\sigma$, $\zeta$ and $\wp$ functions associated with $\Gamma$ and $\mathcal{E}$. Since $\omega\Gamma=\Gamma$, we have
@@
[…]
@@ 2.2 Rational Cherednik algebra for $W=\mathbb{Z}_{m}$
One can view such $y$ as a deformation of $\hbar\frac{d}{dq}$, with the deformation paramters $c_{l}\in\mathbb{C}$ associated with the elements in $\mathbb{Z}_{m}\setminus\{1\}$.
@@
Then one finds that under $\theta$, the elements $e\,q^{m}\,e$, $e\,y^{m}\,e$, and eqyee\,qy\,e are mapped to $u,w$, and $v-\mu_{0}$, respectively. We denote by $B_{\hbar,c}$ the spherical subalgebra in the Dunkl representation:
@@
$v,u]=\hbar mu\,,\quad[w,v]=\hbar mw\,,\quad uw=P_{\hbar}(v)\,,\quad wu=P_{\hbar}(v+\hbar m)\,$, (2.20)
@@ 2.3 Elliptic Cherednik algebra of rank one
We now proceed to define the elliptic version of $H_{\hbar,c}$, following the general framework of [15]. Let $\mathcal{E}=\mathbb{C}/\Gamma$ be an elliptic curve with the symmetry group $\mathbb{Z}_{m}$, $m\in\{2,3,4,6\}$. The elliptic Cherednik algebra $H_{\hbar,c}(\mathcal{E})$ depends on a set of parameters chosen as follows. To every $x_{i}\in\mathcal{E}$ and $l=1,\dots,m-1$ such that $s^{l}(x_{i})=x_{i}$, we assign a parameter $c_{l}(x_{i})$, with an additional requirement that $c_{l}(x_{i})=c_{l}(x_{j})$ whenever $x_{j}=wx_{i}$ for some $w\in W$. From Fig. 1 we observe that this amounts to $4$ parameters in the $m=2$ case, and $6$, 77 or $8$ parameters when $m=3$, 44 or $6$, respectively. We write $c=(c_{l}(x_{i}))$ for the set of parameters. It will be convenient to extend the set $c$ by setting $c_{l}(x_{i})=0$ if $s^{l}(x_{i})\neq x_{i}$. Later it will also be convenient to use the following combinations:
@@
This defines the sheaf $H_{\hbar,c}(\mathcal{E})$ of elliptic Cherednik algebras on $\mathcal{E}$. The sheaf of spherical subalgebras $eH_{\hbar,c}(\mathcal{E})e$ is obtained by replacing local sections $a\in\mathcal{D}_{\mathcal{E}}\rtimes\mathbb{Z}_{m}$ by eaeeae. Again, these local sections can be realised as differential operators using the map (2.16). This produces a sheaf of algebras $B_{\hbar,c}(\mathcal{E}):=\theta\left(eH_{\hbar,c}(\mathcal{E})e\right)\subset\mathcal{D}_{\mathcal{E}}^{\mathbb{Z}_{m}}$.
@@ 2.5 Hamiltonians
To prove this, we notice that any global section $\widehat{h}$ restricts onto $\mathcal{E}_{0}=\mathcal{E}\setminus\{0\}$, and so it must be a combinations of elements $w_{i}$, $v^{(i)}_{j}$. For $\widehat{h}$ to have degree $m$ in $\widehat{p}$, it must be obtained from $w_{m}$ by adding a finite linear combination of $v^{(i)}_{j}$ with $0\leq j\leq m-1$. On the other hand, near $q=0$ each global section must belong to the (completion) of the local rational spherical Cherednik subalgebra, generated by $\mathbb{C}[[u]]$, $v$ and $w$ as given in (2.18). Since $u,v$ are regular at $q=0$, we must have
@@
Comparing this with the previous formula, we conclude that the only allowed terms in $\widehat{h}$ are $v^{(i)}_{j}$ with $j=0,\dots,m-2$ and i+j=m2i+j=m-2, thus proving (2.40).
@@ 2.6 Explicit formulas
Let us write the hamiltonians explicitly, based on the formula (2.40). We use the notation $\omega_{1,2,3}$ and $\eta_{1,2}$ for the fixed points, as in Fig. 1. The parameters $\alpha_{i}$ below are determined through (2.44).
@@ 3.1 Classical dynamics
defines a non-vanishing holomorphic 11-form on $\Sigma$. Hence, the $\mathbb{Z}_{m}$-quotient of the fibration (3.2) defines an elliptic fibration on $T^{*}\mathcal{E}/\mathbb{Z}_{m}$, and the dynamics (3.1) becomes linear along the fibers.
@@ 3.2 Elliptic Dunkl operators
Let us fix some notation. For an abelian variety $X$ with an action of a finite group $W$, the Dunkl operators [19] depend on an auxiliary variable $\alpha\in X^{\vee}=\mathrm{Pic}_{0}(X)$. We will be dealing with the case $X=\mathcal{E}=\mathbb{C}/\Gamma$ and $W=\mathbb{Z}_{m}$, in which case we identify $X^{\vee}\simeq X$ so that $\alpha\in\mathcal{E}$. Below $\sigma$, $\zeta$, $\wp$ stand for the Weierstrass functions associated to the lattice $\Gamma=2\omega_{1}\mathbb{Z}+2\omega_{2}\mathbb{Z}$, and we write $\varphi(x,z)$ for the following combination:
@@
Recall that for each fixed point $x_{i}\in\mathcal{E}$ we have parameters $c_{l}(x_{i})$, $1\leq l\leq m-1$, with $c_{l}(x_{i})=0$ unless $s^{l}(x_{i})=x_{i}$. The fixed points of $s^{l}$ are identified with the cosets
@@
$v_{l}(x,z)$ $=\sum_{\{x_{i}\}}c_{l}(x_{i})e^{-\eta(\Omega_{l}x_{i})z}\varphi(x-x_{i},\Omega_{-l}z)\,,\qquad l\in\mathbb{Z}_{m}\setminus\{0\}\,$, (3.8)
@@
$\eta(\gamma)=-\int_{q}^{q+\gamma}\wp(q)\,dq=\zeta(q+\gamma)-\zeta(q)=2n_{1}\zeta(\omega_{1})+2n_{2}\zeta(\omega_{2})\,.$ (3.9)